OpenStudy (anonymous):

how do you find the perimeter of a triangle when one of the sides is unknown.

OpenStudy (anonymous):

If all you have is the length of two sides of a triangle and nothing else, you can find the last side length by going through the Pythagorean Theorem. the formula for the Pythagorean theorem is a^2 + b^2 = c^2. Say the two lengths of the triangle are 3 and 2. to find the hypotenuse, or missing length, you must plug in the numbers. the two missing sides, or legs, are going to be a and b. so your equation would now look like this: 3^2 + 2^2 = c^2. 3^2 = 9, and 2^2 = 4. 9 +4 = 13. so your equation is: 13= c^2. next you would find the square root of 13 (which you would need a calculator) and the answer would be about 3.6. that would be the last length for your triangle.

OpenStudy (anonymous):

a^2 + b^2 = c^2 then add all the sides up

OpenStudy (anonymous):

If I'm not mistaken, the Pythagorean theorem works just for right triangles right?

OpenStudy (anonymous):

yes juan.

OpenStudy (anonymous):

thanks for reassuring; determining the missing side would then depend on the type of triangle. if it's a right triangle then Pythagorean theory would work. if not, then perhaps you are given angles if so then using trig: (sin, cos, and tan) would give you missing side.

OpenStudy (anonymous):

Thanks!

OpenStudy (precal):

Laws of sine works for all triangles